Crew 182 Operations Report 16Nov2017
Name of person filing report: Carmen Atauconcha
WATER HEATER: Finally, we found the problem. The new water heater is not installed, yet.
CHARGERS OF THE SUITS: Two of the chargers are working, but we moved them and now it is hard for us to connect them, again.
Generator— It worked the whole night since 16:30pm of 15Nov2017 until 12:00 pm of 16Nov2017. Turn on at 16:50 pm 16Nov2017. Operand.
READING FROM 16 NOV 2017
Diesel – 50 %
Propane – 49 psi
Gasoline (5 Gallon containers for ATV) – 2 ½ gallons.
Water (trailer) – 0 gallons
Water (static) – 300 gallons
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – yes
Water Meter: 60 gallons
Toilet tank emptied: yes
ATV’s Used: No
Oil Checked: No
Atv # Fuel Used Gals: 0 gallons
Tires Status: Perfect
# Hours the ATVs were Used today: 0 hours
Deimos used: No
Beginning charge: –
Ending charge: –
Currently charging: Yes
HabCar used? No
Notes and Comments:
I would call today, the day of cleaning. Everyone here is tired. Literally, we cleaned every hole here. There was too much dust. Today in the morning, the commander went to search transportation to return to the Earth. We are sad to leave Mars, and at the same time, happy for returning at home.
From another hand, survivor man is coming. We are so excited¡
Summary of Hab Operations:
We cleaned the whole hab. We found, a couple of glasses and gloves, behind the area of the tools.
Summary of GreenHab operations:
We cleaned the greenhouse. We found too much dust. The two wheelbarrows are in the greenhouse, now.
Summary of ScienceDome operations:
The science dome is finally cleaned. In fact, we found tons of dust under the carpet.
We are using the lamps of the Hab, for the laboratory.
Summary of health and safety issues:
Now, we can say the station is in equilibrium with our health and safety.
Questions, concerns and requests to Mission Support:
We need gasoline for the ATV’S, please. URGENT